Word problems are my toughest challenge. For example;
Draining the vat. With only the small valve open, all of
the liquid can be drained from a large vat in 4 hours. With
only the large valve open, all of the liquid can be drained
from the same vat in 2 hours. How long would it take to
drain the vat with both valves open? How do I figure out the time it takes to drain the vat? Thank you.

Answer by stanbon(48558)

Draining the vat.
With only the small valve open, all of the liquid can be
drained from a large vat in 4 hours.
Time = 4 hrs/job ; Rate = 1/4 job/hr
--------------------------------------------
With only the large valve open, all of the liquid can be drained
from the same vat in 2 hours.
Time = 2 hrs/job ; Rate = 1/2 job/hr.
-----------------------------------------------
How long would it take to drain the vat with both valves open?
Time = x hrs/job ; Rate = 1/x job/hr.
------------------------------------------------
Equation:
rate + rate = together rate
1/4 + 1/2 = 1/x
Multiply thru by 4x to get:
x + 2x = 4
3x = 4
x = 4/3 hr
x = 1 hr and 20 minutes (time to do the job when both vales are open)
==========================================================
Cheers,
Stan H.